OpenStudy (anonymous):

can any1 help me, question posted in comments

OpenStudy (anonymous):

f(x) = sin^4(x+pi/6) + cos^4 (x+pi/6) is a) not a periodic fn b) a periodic fn with period 2pi c) a periodic fn with period pi d) a periodic fn with period pi/2 2. For the hyperbola 9x^2-16y^2-18x+32y-151=0 a) one of the directrix is x =21/5 b) length of latus rectum = 9/2 c) focii are (6,1) & (-4,1) d) ecentricity is 5/4

OpenStudy (nilankshi):

good kush